A squirrel is running westward along the shore at 2.0km/h. The water flows to the east at 7.0km/h relative to the shore. A boat is also traveling to the east at 30km/h relative to the water. If an airplane is traveling east at 150km/h relative to the squirrel, then what is the velocity of the boat to the airplane? Set the eastward direction as positive.
Answer #1

The speed of the boat relative to shore is 30 + 7 = 37 km/h to the east

The speed of the boat relative to the squirrel is 37+2 = 39 km/h to the east

The velocity of the boat relative to the airplane is 39-150 = -111 km/h to the east

or 111 km/h to the west